# Evaluating multiple models using labeled and unlabeled data

### v2

@@ D.1 Results reporting mean absolute error
Tables 1, S3, S4, and S5 report our results on each binary task, for accuracy, ECE, AUC, and AUPRC, respectively. Three high-level findings emerge. First, SSME-KDE achieves the lowest mean absolute error (averaging across tasks and amounts of labeled data). Second, SSME-KDE consistently outperforms the ablated version of SSME, fit to a single model at a time (SSME-KDE-M). And finally, SSME-KDE is able to produce performance estimates that are quite close, in absolute terms, to ground truth. For example, when given 20 labeled examples and 1000 unlabeled examples, SSME-KDE estimates accuracy within at most 2.5 percentage points of ground truth accuracy (across tasks).
@@
Tables S6 and S7 report our results on the multiclass tasks, for accuracy and ECE respectively. Note that we exclude Bayesian-Calibration from multiclass comparisons because the method does not natively support multiclass recalibration. We also omit AutoEval from Table S7 because the implementation of expected calibration error within the framework is not straightforward.
@@
Dataset $n_{\ell}$ $n_{u}$ Labeled Majority-Vote Pseudo-Labeling (LR) Dawid-Skene Bayesian-Calibration SSME-KDE-M SSME-KDE (Ours)
Critical Outcome 20 1000 11.01 ± 4.04 4.49 ± 0.31 6.94 ± 2.30 2.61 ± 0.33 3.48 ± 2.76 3.17 ± 1.10 1.16 ± 0.48
50 1000 6.22 ± 2.23 4.30 ± 0.41 5.38 ± 1.40 2.37 ± 0.32 2.56 ± 1.57 3.01 ± 0.94 1.13 ± 0.47
100 1000 4.20 ± 1.38 4.05 ± 0.43 3.63 ± 0.77 2.25 ± 0.37 1.69 ± 0.91 2.81 ± 0.81 1.15 ± 0.38
ED Revisit 20 1000 8.37 ± 3.14 1.83 ± 0.07 4.16 ± 2.96 3.25 ± 2.45 3.57 ± 2.78 1.88 ± 0.86 0.76 ± 0.16
50 1000 4.82 ± 1.73 1.78 ± 0.11 2.29 ± 1.69 2.29 ± 1.68 2.04 ± 1.66 1.83 ± 0.67 0.73 ± 0.18
100 1000 3.29 ± 0.88 1.70 ± 0.14 1.36 ± 0.78 1.34 ± 0.76 1.16 ± 0.73 1.51 ± 0.63 0.73 ± 0.21
Hospital Admission 20 1000 21.76 ± 4.18 17.34 ± 0.47 8.10 ± 4.61 17.31 ± 0.42 5.12 ± 3.94 5.54 ± 1.32 1.97 ± 0.47
50 1000 12.74 ± 2.25 16.63 ± 0.48 5.02 ± 2.45 16.60 ± 0.43 3.49 ± 2.05 5.20 ± 1.19 2.06 ± 0.67
100 1000 8.56 ± 1.39 15.71 ± 0.46 3.91 ± 1.76 15.62 ± 0.44 3.23 ± 1.68 5.32 ± 1.32 1.70 ± 0.54
SARS-CoV Inhibition 20 1000 7.44 ± 3.44 4.57 ± 0.18 5.96 ± 3.13 4.35 ± 0.53 2.24 ± 1.19 2.57 ± 0.64 3.38 ± 0.47
50 1000 3.66 ± 1.80 4.59 ± 0.16 3.06 ± 1.28 4.08 ± 0.57 1.73 ± 0.93 2.27 ± 0.72 3.41 ± 0.41
100 1000 2.18 ± 1.14 4.65 ± 0.19 2.36 ± 0.78 3.67 ± 0.59 1.35 ± 0.78 1.79 ± 0.69 3.44 ± 0.47
Toxicity Detection 20 1000 5.85 ± 2.89 4.15 ± 0.32 5.09 ± 2.87 4.40 ± 0.33 4.69 ± 1.21 5.67 ± 0.68 2.35 ± 0.46
50 1000 3.99 ± 2.28 3.99 ± 0.29 3.04 ± 1.66 4.20 ± 0.26 3.97 ± 1.21 4.57 ± 0.93 2.26 ± 0.44
100 1000 2.37 ± 1.35 3.86 ± 0.38 1.91 ± 0.99 4.10 ± 0.32 3.30 ± 0.91 3.43 ± 1.05 2.19 ± 0.53
@@
Dataset $n_{\ell}$ $n_{u}$ Labeled Majority-Vote Pseudo-Labeling (LR) Dawid-Skene Bayesian-Calibration SSME-KDE-M SSME-KDE (Ours)
Critical Outcome 20 1000 10.09 ± 4.84 9.80 ± 1.15 31.73 ± 3.95 9.39 ± 1.25 2.84 ± 0.91 4.72 ± 2.27 2.52 ± 1.24
50 1000 7.50 ± 4.62 8.19 ± 2.12 27.33 ± 5.51 8.49 ± 1.46 3.17 ± 1.17 5.61 ± 4.61 2.39 ± 1.74
100 1000 5.65 ± 3.44 7.31 ± 2.09 20.43 ± 4.38 7.97 ± 1.08 2.70 ± 0.94 3.82 ± 1.72 2.83 ± 2.89
ED Revisit 20 1000 18.48 ± 6.68 19.48 ± 8.75 7.48 ± 0.72 8.27 ± 3.80 7.65 ± 0.55 11.89 ± 4.66 5.92 ± 3.14
50 1000 17.37 ± 7.13 17.97 ± 9.26 7.48 ± 0.95 7.62 ± 0.99 7.30 ± 0.76 11.99 ± 4.36 5.09 ± 2.56
100 1000 14.13 ± 6.03 14.84 ± 8.40 7.06 ± 1.46 7.09 ± 1.52 7.47 ± 1.17 11.28 ± 5.73 5.08 ± 2.77
Hospital Admission 20 1000 6.97 ± 4.64 16.66 ± 0.30 8.94 ± 5.97 16.70 ± 0.31 2.67 ± 1.15 3.63 ± 1.95 2.51 ± 1.38
50 1000 5.08 ± 3.49 16.09 ± 0.36 5.59 ± 4.31 16.18 ± 0.31 2.62 ± 1.65 3.18 ± 1.95 2.51 ± 1.20
100 1000 3.57 ± 2.58 15.29 ± 0.43 3.66 ± 2.68 15.32 ± 0.39 2.55 ± 1.34 3.17 ± 1.60 2.02 ± 1.20
SARS-CoV Inhibition 20 1000 9.61 ± 9.22 7.64 ± 1.21 30.92 ± 4.35 7.50 ± 1.05 3.07 ± 1.00 5.42 ± 2.63 3.48 ± 1.58
50 1000 5.84 ± 3.64 6.86 ± 1.59 22.71 ± 4.29 7.06 ± 1.04 3.62 ± 0.97 5.02 ± 1.86 3.41 ± 1.68
100 1000 3.97 ± 1.97 5.33 ± 1.60 16.33 ± 3.27 6.04 ± 1.17 3.53 ± 1.35 4.21 ± 1.90 3.46 ± 1.63
Toxicity Detection 20 1000 6.71 ± 3.57 6.54 ± 0.32 17.32 ± 7.51 6.20 ± 0.41 5.22 ± 0.59 6.05 ± 1.02 3.34 ± 0.82
50 1000 4.76 ± 3.29 6.30 ± 0.28 11.79 ± 6.41 5.97 ± 0.33 4.76 ± 0.74 4.86 ± 1.03 3.15 ± 0.66
100 1000 3.82 ± 2.17 6.14 ± 0.44 7.54 ± 3.73 5.84 ± 0.44 4.25 ± 0.96 4.15 ± 1.20 3.09 ± 0.81
@@
Dataset $n_{\ell}$ $n_{u}$ Labeled Majority-Vote Pseudo-Labeling (LR) Dawid-Skene Bayesian-Calibration SSME-KDE-M SSME-KDE (Ours)
Critical Outcome 20 1000 32.86 ± 18.26 22.41 ± 8.36 22.98 ± 6.69 39.02 ± 4.26 9.29 ± 6.01 11.48 ± 5.46 6.11 ± 2.63
50 1000 22.81 ± 13.16 17.44 ± 7.60 20.48 ± 8.04 35.64 ± 5.80 9.34 ± 5.17 11.98 ± 5.17 6.17 ± 3.47
100 1000 15.71 ± 8.81 16.22 ± 6.66 14.45 ± 7.30 33.31 ± 4.33 8.96 ± 5.07 11.30 ± 6.01 5.77 ± 2.80
ED Revisit 20 1000 19.18 ± 13.27 2.63 ± 0.44 5.14 ± 3.20 5.12 ± 4.68 9.14 ± 3.74 5.07 ± 2.89 1.67 ± 1.06
50 1000 8.85 ± 8.14 2.44 ± 0.34 2.72 ± 2.22 3.04 ± 2.80 6.03 ± 2.95 3.79 ± 2.33 1.81 ± 1.08
100 1000 6.34 ± 5.57 2.46 ± 1.73 1.57 ± 1.19 1.74 ± 1.24 4.23 ± 1.97 3.92 ± 2.23 1.82 ± 1.13
Hospital Admission 20 1000 9.43 ± 5.85 20.87 ± 0.51 10.89 ± 9.33 21.15 ± 0.52 5.26 ± 3.84 4.36 ± 1.60 3.47 ± 2.04
50 1000 7.46 ± 4.74 19.96 ± 0.70 7.91 ± 5.89 20.34 ± 0.59 4.43 ± 2.68 3.70 ± 2.26 3.64 ± 2.16
100 1000 5.51 ± 3.48 18.97 ± 0.70 4.12 ± 3.66 19.17 ± 0.68 3.49 ± 2.17 4.00 ± 2.19 2.80 ± 1.84
SARS-CoV Inhibition 20 1000 22.27 ± 10.94 17.50 ± 3.17 37.41 ± 8.82 16.60 ± 3.91 7.54 ± 2.74 13.81 ± 5.52 20.51 ± 6.12
50 1000 15.02 ± 8.77 14.94 ± 4.23 30.29 ± 9.40 15.01 ± 3.85 8.40 ± 3.45 12.82 ± 3.63 21.06 ± 5.46
100 1000 11.53 ± 5.64 13.37 ± 2.75 20.34 ± 6.46 12.61 ± 3.78 8.27 ± 3.19 11.01 ± 5.02 20.67 ± 5.92
Toxicity Detection 20 1000 19.34 ± 8.45 24.12 ± 1.49 25.12 ± 12.69 26.34 ± 1.31 19.94 ± 4.70 23.38 ± 2.64 10.89 ± 3.14
50 1000 13.78 ± 6.52 23.11 ± 1.43 20.15 ± 12.57 25.24 ± 1.31 16.84 ± 5.68 18.90 ± 3.88 9.91 ± 3.06
100 1000 10.69 ± 6.16 22.34 ± 1.99 14.06 ± 7.21 24.51 ± 1.68 14.15 ± 5.35 14.59 ± 4.54 9.88 ± 3.51

### v3

@@ D.2 Results reporting mean absolute error
Tables S3, S4, S5, and S6 report results on each binary task for accuracy, ECE, AUC, and AUPRC, respectively. Three high-level findings emerge. First, SSME-KDE achieves the lowest mean absolute error (averaging across tasks and amounts of labeled data). Second, SSME-KDE consistently outperforms the ablated version of SSME, fit to a single model at a time (SSME-KDE-M). And finally, SSME-KDE is able to produce performance estimates that are quite close, in absolute terms, to ground truth. For example, when given 20 labeled examples and 1000 unlabeled examples, SSME-KDE estimates accuracy within at most 2.5 percentage points of ground truth accuracy (across tasks).
@@
Tables S11 and S12 report our results on the multiclass tasks, for accuracy and ECE respectively. Note that we exclude Bayesian-Calibration from multiclass comparisons because the method does not natively support multiclass recalibration. We also omit AutoEval from Table S12 because the implementation of expected calibration error within the framework is not straightforward.
@@
Dataset $n_{\ell}$ $n_{u}$ Labeled Majority-Vote PL Dawid-Skene SPE BC AutoEval Active-Testing SSME-M SSME (Ours)
Critical Outcome 20 1000 5.19 ± 1.07 4.01 ± 0.13 4.12 ± 1.07 4.36 ± 0.09 12.73 ± 5.64 2.80 ± 0.62 4.78 ± 0.93 5.17 ± 0.32 1.70 ± 0.27 0.67 ± 0.13
50 1000 2.90 ± 0.59 4.21 ± 0.12 3.06 ± 0.64 4.07 ± 0.11 16.79 ± 9.66 2.07 ± 0.36 3.01 ± 0.65 5.61 ± 0.40 1.65 ± 0.25 0.78 ± 0.13
100 1000 2.09 ± 0.41 4.10 ± 0.14 1.58 ± 0.30 3.87 ± 0.11 8.61 ± 8.84 1.18 ± 0.21 2.00 ± 0.32 5.48 ± 0.35 1.30 ± 0.20 0.77 ± 0.13
ED Revisit 20 1000 5.11 ± 0.98 2.15 ± 0.02 5.13 ± 0.89 4.02 ± 0.78 14.70 ± 10.59 4.36 ± 0.77 4.70 ± 0.92 2.83 ± 0.20 1.64 ± 0.34 0.45 ± 0.10
50 1000 2.02 ± 0.58 2.10 ± 0.03 2.73 ± 0.62 2.74 ± 0.61 17.96 ± 7.41 2.47 ± 0.57 1.95 ± 0.57 3.03 ± 0.26 1.46 ± 0.27 0.53 ± 0.11
100 1000 1.43 ± 0.32 2.00 ± 0.05 1.54 ± 0.34 1.51 ± 0.33 11.69 ± 5.13 1.43 ± 0.31 1.42 ± 0.29 2.64 ± 0.16 1.18 ± 0.25 0.57 ± 0.11
Hospital Admission 20 1000 7.32 ± 1.25 16.32 ± 0.66 6.86 ± 1.19 19.55 ± 0.15 14.27 ± 1.76 2.48 ± 0.44 7.19 ± 1.03 9.33 ± 0.84 3.29 ± 0.47 1.88 ± 0.29
50 1000 5.40 ± 0.83 16.37 ± 0.62 3.99 ± 0.82 18.78 ± 0.16 15.68 ± 0.69 2.14 ± 0.36 5.23 ± 0.68 9.25 ± 0.64 3.17 ± 0.51 1.95 ± 0.27
100 1000 3.64 ± 0.55 15.34 ± 0.66 3.01 ± 0.53 17.81 ± 0.18 14.32 ± 0.95 2.42 ± 0.33 4.01 ± 0.55 9.24 ± 0.80 3.06 ± 0.45 1.51 ± 0.23
SARS-CoV Inhibition 20 1000 6.11 ± 0.96 2.29 ± 0.42 5.95 ± 1.00 4.91 ± 0.18 11.42 ± 5.06 2.25 ± 0.31 4.59 ± 0.84 5.97 ± 0.47 3.06 ± 0.23 2.30 ± 0.16
50 1000 3.22 ± 0.57 2.46 ± 0.38 2.99 ± 0.45 4.50 ± 0.17 8.04 ± 2.56 1.74 ± 0.21 2.64 ± 0.42 5.74 ± 0.32 2.59 ± 0.26 2.35 ± 0.10
100 1000 2.04 ± 0.38 2.38 ± 0.35 2.14 ± 0.31 4.01 ± 0.17 6.09 ± 4.63 1.43 ± 0.19 1.94 ± 0.25 5.99 ± 0.47 1.84 ± 0.24 2.36 ± 0.13
Toxicity Detection 20 1000 5.95 ± 0.73 4.97 ± 0.31 5.03 ± 0.81 4.82 ± 0.09 11.26 ± 5.18 5.29 ± 0.29 5.27 ± 0.75 7.19 ± 0.44 6.71 ± 0.23 2.34 ± 0.15
50 1000 4.03 ± 0.68 4.79 ± 0.26 2.88 ± 0.48 4.65 ± 0.08 8.75 ± 1.66 4.57 ± 0.30 3.37 ± 0.41 7.26 ± 0.47 5.38 ± 0.28 2.22 ± 0.13
100 1000 2.43 ± 0.41 4.46 ± 0.22 1.90 ± 0.31 4.46 ± 0.11 7.41 ± 0.16 3.78 ± 0.26 2.34 ± 0.26 7.50 ± 0.60 3.80 ± 0.32 2.14 ± 0.15
Table S3: Mean absolute error in accuracy estimation on binary tasks. We report mean absolute error (MAE, averaging across classifiers) across five binary classification tasks and different amounts of labeled data. Each entry corresponds to the mean MAE across 50 randomized splits of data, accompanied by the 95% CI (computed as 1.96 $\times$ the standard error) in MAE across splits. We bold the best performing method in each row, and underline the next best performing method by mean MAE. SSME-M refers to performance when fitting SSME to a single classifier’s scores, instead of modeling the joint distribution of $P(y,\mathbf{s})$.
Dataset $n_{\ell}$ $n_{u}$ Labeled Majority-Vote PL Dawid-Skene SPE BC SSME-M SSME (Ours)
Critical Outcome 20 1000 11.01 ± 1.12 2.75 ± 0.26 6.94 ± 0.64 2.61 ± 0.09 16.17 ± 6.69 3.48 ± 0.76 3.17 ± 0.30 1.16 ± 0.13
50 1000 6.22 ± 0.62 3.68 ± 0.34 5.38 ± 0.39 2.37 ± 0.09 20.91 ± 9.43 2.56 ± 0.44 3.01 ± 0.26 1.13 ± 0.13
100 1000 4.20 ± 0.38 3.43 ± 0.28 3.63 ± 0.21 2.25 ± 0.10 12.84 ± 8.93 1.69 ± 0.25 2.81 ± 0.22 1.15 ± 0.10
ED Revisit 20 1000 8.37 ± 0.87 1.81 ± 0.02 4.16 ± 0.82 3.25 ± 0.68 14.15 ± 10.31 3.57 ± 0.77 1.88 ± 0.24 0.76 ± 0.04
50 1000 4.82 ± 0.48 1.77 ± 0.03 2.29 ± 0.47 2.29 ± 0.46 17.12 ± 7.09 2.04 ± 0.46 1.83 ± 0.19 0.73 ± 0.05
100 1000 3.29 ± 0.24 1.69 ± 0.04 1.36 ± 0.22 1.34 ± 0.21 10.85 ± 4.82 1.16 ± 0.20 1.51 ± 0.17 0.73 ± 0.06
Hospital Admission 20 1000 21.76 ± 1.16 21.73 ± 2.75 8.10 ± 1.28 17.31 ± 0.13 16.79 ± 0.58 5.12 ± 1.09 5.54 ± 0.36 1.97 ± 0.13
50 1000 12.74 ± 0.62 21.03 ± 2.65 5.02 ± 0.68 16.60 ± 0.13 14.87 ± 0.47 3.49 ± 0.57 5.20 ± 0.33 2.06 ± 0.18
100 1000 8.56 ± 0.38 19.97 ± 2.45 3.91 ± 0.49 15.62 ± 0.13 13.39 ± 0.79 3.23 ± 0.47 5.32 ± 0.37 1.70 ± 0.15
SARS-CoV Inhibition 20 1000 7.44 ± 0.95 1.74 ± 0.38 5.96 ± 0.87 4.35 ± 0.15 11.85 ± 5.63 2.24 ± 0.33 2.57 ± 0.18 3.38 ± 0.13
50 1000 3.66 ± 0.50 1.96 ± 0.39 3.06 ± 0.35 4.08 ± 0.16 8.96 ± 2.74 1.73 ± 0.26 2.27 ± 0.20 3.41 ± 0.11
100 1000 2.18 ± 0.32 1.83 ± 0.37 2.36 ± 0.22 3.67 ± 0.16 7.28 ± 4.71 1.35 ± 0.22 1.79 ± 0.19 3.44 ± 0.13
Toxicity Detection 20 1000 5.85 ± 0.80 5.08 ± 0.39 5.09 ± 0.79 4.40 ± 0.09 11.23 ± 5.14 4.69 ± 0.34 5.67 ± 0.19 2.35 ± 0.13
50 1000 3.99 ± 0.63 4.84 ± 0.33 3.04 ± 0.46 4.20 ± 0.07 7.97 ± 1.76 3.97 ± 0.34 4.57 ± 0.26 2.26 ± 0.12
100 1000 2.37 ± 0.37 4.63 ± 0.33 1.91 ± 0.27 4.10 ± 0.09 6.75 ± 0.15 3.30 ± 0.25 3.43 ± 0.29 2.19 ± 0.15
@@
Dataset $n_{\ell}$ $n_{u}$ Labeled Majority-Vote PL Dawid-Skene SPE BC SSME-M SSME (Ours)
Critical Outcome 20 1000 10.09 ± 1.34 9.45 ± 0.39 31.73 ± 1.10 9.39 ± 0.35 10.92 ± 4.41 2.84 ± 0.25 4.72 ± 0.63 2.52 ± 0.34
50 1000 7.50 ± 1.28 8.06 ± 0.53 27.33 ± 1.53 8.49 ± 0.40 20.24 ± 11.78 3.17 ± 0.32 5.61 ± 1.28 2.39 ± 0.48
100 1000 5.65 ± 0.95 7.29 ± 0.61 20.43 ± 1.22 7.97 ± 0.30 13.71 ± 10.26 2.70 ± 0.26 3.82 ± 0.48 2.83 ± 0.80
ED Revisit 20 1000 18.48 ± 1.85 19.99 ± 2.44 7.48 ± 0.20 8.27 ± 1.05 25.84 ± 6.42 7.65 ± 0.15 11.89 ± 1.29 5.92 ± 0.87
50 1000 17.37 ± 1.98 17.93 ± 2.58 7.48 ± 0.26 7.62 ± 0.28 29.83 ± 4.67 7.30 ± 0.21 11.99 ± 1.21 5.09 ± 0.71
100 1000 14.13 ± 1.67 14.95 ± 2.29 7.06 ± 0.40 7.09 ± 0.42 27.00 ± 3.59 7.47 ± 0.32 11.28 ± 1.59 5.08 ± 0.77
Hospital Admission 20 1000 6.97 ± 1.29 16.20 ± 1.00 8.94 ± 1.65 16.70 ± 0.10 15.81 ± 0.33 2.67 ± 0.32 3.63 ± 0.54 2.51 ± 0.38
50 1000 5.08 ± 0.97 15.47 ± 0.56 5.59 ± 1.20 16.18 ± 0.10 14.93 ± 0.40 2.62 ± 0.46 3.18 ± 0.54 2.51 ± 0.33
100 1000 3.57 ± 0.71 14.78 ± 0.44 3.66 ± 0.74 15.32 ± 0.12 13.95 ± 0.72 2.55 ± 0.37 3.17 ± 0.44 2.02 ± 0.33
SARS-CoV Inhibition 20 1000 9.61 ± 2.56 5.44 ± 0.60 30.92 ± 1.21 7.50 ± 0.29 13.72 ± 5.03 3.07 ± 0.28 5.42 ± 0.73 3.48 ± 0.44
50 1000 5.84 ± 1.01 5.22 ± 0.39 22.71 ± 1.19 7.06 ± 0.29 10.67 ± 2.28 3.62 ± 0.27 5.02 ± 0.52 3.41 ± 0.47
100 1000 3.97 ± 0.55 4.95 ± 0.35 16.33 ± 0.91 6.04 ± 0.32 10.34 ± 4.18 3.53 ± 0.38 4.21 ± 0.53 3.46 ± 0.45
Toxicity Detection 20 1000 6.71 ± 0.99 5.45 ± 0.82 17.32 ± 2.08 6.20 ± 0.11 12.38 ± 5.76 5.22 ± 0.16 6.05 ± 0.28 3.34 ± 0.23
50 1000 4.76 ± 0.91 5.15 ± 0.47 11.79 ± 1.78 5.97 ± 0.09 8.23 ± 1.65 4.76 ± 0.21 4.86 ± 0.29 3.15 ± 0.18
100 1000 3.82 ± 0.60 4.85 ± 0.20 7.54 ± 1.03 5.84 ± 0.12 7.28 ± 0.22 4.25 ± 0.27 4.15 ± 0.33 3.09 ± 0.22
@@
Dataset $n_{\ell}$ $n_{u}$ Labeled Majority-Vote PL Dawid-Skene SPE BC SSME-M SSME (Ours)
Critical Outcome 20 1000 32.86 ± 5.06 36.21 ± 2.09 22.98 ± 1.85 39.02 ± 1.18 45.09 ± 6.81 9.29 ± 1.67 11.48 ± 1.51 6.11 ± 0.73
50 1000 22.81 ± 3.65 26.12 ± 2.77 20.48 ± 2.23 35.64 ± 1.61 44.67 ± 7.31 9.34 ± 1.43 11.98 ± 1.43 6.17 ± 0.96
100 1000 15.71 ± 2.44 24.01 ± 2.47 14.45 ± 2.02 33.31 ± 1.20 47.35 ± 6.71 8.96 ± 1.41 11.30 ± 1.67 5.77 ± 0.77
ED Revisit 20 1000 19.18 ± 3.68 4.53 ± 1.59 5.14 ± 0.89 5.12 ± 1.30 16.66 ± 11.60 9.14 ± 1.04 5.07 ± 0.80 1.67 ± 0.29
50 1000 8.85 ± 2.26 3.12 ± 0.79 2.72 ± 0.62 3.04 ± 0.78 24.55 ± 10.06 6.03 ± 0.82 3.79 ± 0.65 1.81 ± 0.30
100 1000 6.34 ± 1.54 3.23 ± 0.82 1.57 ± 0.33 1.74 ± 0.34 24.75 ± 12.17 4.23 ± 0.55 3.92 ± 0.62 1.82 ± 0.31
Hospital Admission 20 1000 9.43 ± 1.62 32.72 ± 6.41 10.89 ± 2.59 21.15 ± 0.16 21.13 ± 0.34 5.26 ± 1.07 4.36 ± 0.44 3.47 ± 0.57
50 1000 7.46 ± 1.31 31.37 ± 6.28 7.91 ± 1.63 20.34 ± 0.18 19.21 ± 0.80 4.43 ± 0.74 3.70 ± 0.63 3.64 ± 0.60
100 1000 5.51 ± 0.97 29.71 ± 5.98 4.12 ± 1.02 19.17 ± 0.21 18.73 ± 0.71 3.49 ± 0.60 4.00 ± 0.61 2.80 ± 0.51
SARS-CoV Inhibition 20 1000 22.27 ± 3.03 18.75 ± 2.82 37.41 ± 2.44 16.60 ± 1.08 41.22 ± 6.90 7.54 ± 0.76 13.81 ± 1.53 20.51 ± 1.70
50 1000 15.02 ± 2.43 14.10 ± 1.05 30.29 ± 2.61 15.01 ± 1.07 34.97 ± 1.49 8.40 ± 0.96 12.82 ± 1.01 21.06 ± 1.51
100 1000 11.53 ± 1.56 15.72 ± 2.66 20.34 ± 1.79 12.61 ± 1.05 33.60 ± 2.49 8.27 ± 0.89 11.01 ± 1.39 20.67 ± 1.64
Toxicity Detection 20 1000 19.34 ± 2.34 29.05 ± 3.31 25.12 ± 3.52 26.34 ± 0.36 37.10 ± 6.70 19.94 ± 1.30 23.38 ± 0.73 10.89 ± 0.87
50 1000 13.78 ± 1.81 30.54 ± 3.70 20.15 ± 3.48 25.24 ± 0.36 31.94 ± 1.07 16.84 ± 1.58 18.90 ± 1.08 9.91 ± 0.85
100 1000 10.69 ± 1.71 23.87 ± 0.84 14.06 ± 2.00 24.51 ± 0.47 30.94 ± 0.58 14.15 ± 1.48 14.59 ± 1.26 9.88 ± 0.97
@@
D.3 Results reporting absolute performance estimates
Results thus far have reported aggregate errors in performance estimates across classifiers in the set. Here, we include results on a per-classifier basis in the context of toxicity detection, the task for which we have the largest variability in classifier quality (Tables S7, S8, S9, S10). The tables illustrate how SSME’s performance manifests on a per-classifier basis, often producing more accurate estimates than the baselines on the classifiers with lowest performance. The tables also make evident that SSME’s improvement in performance estimation can be attributed to a significant reduction in the variance of performance estimates across different data splits.